Mr Qadri bought a piece of land for RS 36,000 at 120 sq.m . Find the length of the plot if it's breadth is 3/4 of the length
Answer by CubeyThePenguin(3113)

width = w
length = 3w/4

w(3w/4) = 120
3w^2/4 = 120
3w^2 = 480
w^2 = 160

w = 4sqrt(10)
length = 3w/4 = 3sqrt(10)